### v1

# : A Million-scale Real Prompt-Gallery Dataset for Text-to-Video Diffusion Models

Abstract
The arrival of Sora marks a new era for text-to-video diffusion models, bringing significant advancements in video generation and potential applications. However, Sora, as well as other text-to-video diffusion models, highly relies on the prompts, and there is no publicly available dataset featuring a study of text-to-video prompts. In this paper, we introduce VidProM, the first large-scale dataset comprising 1.67 Million unique text-to-Video Prompts from real users. Additionally, the dataset includes 6.69 million videos generated by four state-of-the-art diffusion models and some related data. We initially demonstrate the curation of this large-scale dataset, which is a time-consuming and costly process. Subsequently, we show how the proposed VidProM differs from DiffusionDB, a large-scale prompt-gallery dataset for image generation. Based on the analysis of these prompts, we identify the necessity for a new prompt dataset specifically designed for text-to-video generation and gain insights into the preferences of real users when creating videos. Our large-scale and diverse dataset also inspires many exciting new research areas. For instance, to develop better, more efficient, and safer text-to-video diffusion models, we suggest exploring text-to-video prompt engineering, efficient video generation, and video copy detection for diffusion models. We make the collected dataset VidProM publicly available at GitHub and Hugging Face under the CC-BY-NC 4.0 License.
1 Introduction
The Sora [6] initializes a new era for text-to-video diffusion models, revolutionizing video generation with significant advancements. This breakthrough provides new possibilities for storytelling, immersive experiences, and content creation, as Sora [6] transforms textual descriptions into high-quality videos with ease. However, Sora [6] and other text-to-video diffusion models [9, 1, 3, 5] heavily relies on the prompts used. Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which hinders the development and evaluation of these models.
In this paper, we present the first systematic research on the text-to-video prompts. Specifically, our efforts primarily focus on building the first text-to-video prompt-gallery dataset VidProM, conducting an in-depth comparison with DiffusionDB [22], analyzing prefered topics by users, and introducing new research directions based on our VidProM. The demonstration of VidProM is shown in Fig. 1.
The first text-to-video prompt-gallery dataset. Our large-scale VidProM includes million unique text-to-video prompts from real users and million generated videos by state-of-the-art diffusion models. The prompts are from official Pika Discord channels, and the videos are generated by Pika [9], Text2Video-Zero [1], VideoCraft2 [3], and ModelScope [5]. We distribute the generation process across 10 servers, each equipped with 8 Nvidia V100 GPUs. Each prompt is embedded using the powerful text-embedding-3-large model from OpenAI and assigned six not-safe-for-work (NSFW) probabilities, consisting of toxicity, obscenity, identity attack, insult, threat, and sexual explicitness. We also add a Universally Unique Identifier (UUID) and a time stamp to each data point in our VidProM. In addition to the main dataset, we introduce a subset named VidProS, which consists of semantically unique prompts. In this subset, the cosine similarity between any two prompts is less than 0.8, ensuring a high level of semantic diversity.
In-depth comparison with DiffusionDB and people preference analysis. We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [22]. By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [22] lies in: (1) Our VidProM contains more semantically unique prompts, which are embedded by a more advanced model and collected over a longer time span. (2) We collect videos through both web scraping and local generation, whereas DiffusionDB [22] only contains web-scraped images, resulting in our dataset being more time-consuming and costly to assemble. (3) The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer. These differences highlight the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models. Based on our analysis of the new prompt dataset, we conclude that (1) frequently occurring words include ‘modern’, ‘motion’, ‘close’, ‘forest’, and ‘sky’; and (2) preferred topics encompass humans, science fiction, and animals.
Inspiring new research directions. The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up numerous exciting research directions. Researchers are focused on developing better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering. (2) For more efficient models, researchers can search for related prompts in our VidProM and reconstruct new videos from similar existing videos, thereby avoiding the need to generate videos from scratch. (3) For safer models, researchers can develop specialized models to distinguish generated videos from real videos to combat misinformation, and train video copy detection models to identify potential copyright issues. Apart from diffusion models, text-video pairs are also utilized in multimodal learning tasks such as video-text retrieval and video captioning. Our prompts and synthetic videos can help mitigate copyright concerns associated with using online videos and alleviate the difficulty of collecting high-quality video-text data.
To sum up, this paper makes the following contributions: (1) We contribute the first text-to-video prompt-gallery dataset, VidProM, which includes million unique prompts from real users and million generated videos by state-of-the-art diffusion models. (2) We provide a detailed and in-depth comparison with the text-to-image prompt-gallery dataset, DiffusionDB, and highlight the necessity of VidProM as well as real users’ preference. (3) We reveal several exciting research directions inspired by VidProM and position it as a rich database for future studies.
2 Related Works
2.1 Text-to-Video Diffusion Models
Text-to-Video diffusion models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] have become a powerful tool for producing high-quality video content from textual prompts. Pika [9] is a commercial text-to-video model by Pika Labs, which advances the field of video generation. Text2Video-Zero [1] enables zero-shot video generation using textual prompts. VideoCrafter2 [3] generates videos with high visual quality and precise text-video alignment without requiring high-quality videos. ModelScope [5] evolves from a text-to-image model by adding spatio-temporal blocks for consistent frame generation and smooth movement transitions. This paper uses these four publicly accessible sources (access to generated videos or pre-trained weights) for constructing our VidProM. We hope that the collection of diffusion-generated videos will be useful for further research in text-to-video generation community.
2.2 Existing Datasets
Text-Video Datasets. While several published text-video datasets exist [12, 13, 14, 15, 16, 17, 18, 19], they primarily consist of caption-(real)-video pairs rather than prompt-(generated)-video pairs. For example, WebVid-10M is a large-scale text-video dataset with 10 million video-text pairs collected from stock footage websites [15]. HDVILA-100M [16] is a comprehensive video-language dataset designed for multimodal representation learning, offering high-resolution and diverse content. Panda-70M [17] is a curated subset of HDVILA-100M [16], featuring semantically consistent and high-resolution videos. In contrast, our VidProM contains prompts authored by real users to generate videos of interest, and the videos are produced by text-to-video diffusion models.
Prompt Datasets. Existing datasets underscore the significance of compiling a set of prompts. In the text-to-text domain, studies [20] demonstrate that gathering and analyzing prompts can aid in developing language models that respond more effectively to prompts. PromptSource [21] recognizes the growing popularity of using prompts to train and query language models, and thus create a system for generating, sharing, and utilizing natural language prompts. In the text-to-image domain, DiffusionDB [22] collects a large-scale prompt-image dataset, revealing its potential to open up new avenues for research. Given the importance of prompt datasets and the new era of text-to-video generation brought by Sora [6], this paper presents the first prompt dataset specifically collected for text-to-video generation.
3 Curating VidProM
This section shows the steps of curating our VidProM, including (1) collecting source HTML files, (2) extracting and embedding prompts, (3) assigning NSFW probabilities, and (4) scraping and generating videos. In Fig. 2, we illustrate a single data point in the proposed VidProM. This data point includes a prompt, a UUID, a timestamp, six NSFW probabilities, a 3072-dimensional prompt embedding, and four generated videos. After building VidProM, we filter out prompts with high semantic similarity to each other, resulting in VidProS.
Collecting Source HTML Files. We gather chat messages from the official Pika Discord channels between July 2023 and February 2024 using DiscordChatExporter [23] and store them as HTML files. Our focus is on 10 channels where users input prompts and request a bot to execute the Pika text-to-video diffusion model for video generation. The user inputs and outputs are made available by Pika Lab under the Creative Commons Noncommercial 4.0 Attribution International License (CC BY-NC 4.0), as detailed in Section 4.5.a of their official terms of service. Consequently, the text-to-video prompts and Pika videos in our dataset are open-sourced under the same license.
Extracting and Embedding Prompts. The HTML files are then processed using regular expressions to extract prompts and time stamps. We subsequently filter out prompts used for image-to-video generation (because the images are not publicly available) and prompts without associated videos (these prompts may have been banned by Pika or hidden by the users). Finally, we remove duplicate prompts and assign a UUID to each one, resulting in a total of unique prompts. Because the text-to-video prompts are significantly complex and long, we use OpenAI’s text-embedding-3-large API, which supports up to tokens, to embed all of our prompts. We retain the original 3072-dimensional output of the text-embedding-3-large model, allowing users of our dataset to freely reduce their dimensionality according to their preferences using Principal Component Analysis (PCA).
Assigning NSFW Probabilities. We select the public Discord channels of Pika Labs, which prohibit NSFW content, as the source for our text-to-video prompts. Consequently, if a user submits a harmful prompt, the channel will automatically reject it. However, we find VidProM still includes NSFW prompts that were not filtered by Pika. We employ a state-of-the-art NSFW model, Detoxify [24], to assign probabilities in six aspects of NSFW content, including toxicity, obscenity, identity attack, insult, threat, and sexual explicitness, to each prompt. In Fig. 3, we visualize the number of prompts with a NSFW probability greater than . We conclude that only a very small fraction (less than ) of prompts have a probability greater than . Our dataset, VidProM, provides six separate NSFW probabilities, allowing researchers to determine a suitable threshold for filtering out potentially unsafe data specific to their tasks.
Scraping and Generating Videos. We enhance our VidProM diversity by not only scraping Pika videos from extracted links but also utilizing three state-of-the-art open-source text-to-video diffusion models for video generation. This process demands significant computational resources: we distribute the text-to-video generation across 10 servers, each equipped with 8 Nvidia V100 GPUs. It costs us approximately 50,631 GPU hours and results in 6.69 million videos (), totaling 24,414,747.8 seconds in duration. The breakdown of video lengths is as follows: 3.0 seconds for Pika, 2.0 seconds for Text2Video-Zero, 1.6 seconds for VideoCraft2, and 8.0 seconds for ModelScope.
Selecting Semantically Unique Prompts. Beyond general uniqueness, we introduce a new concept: semantically unique prompts. We define a dataset as containing only semantically unique prompts if, for any two arbitrary prompts, their cosine similarity calculated using text-embedding-3-large embeddings is less than . After semantic de-duplication, our VidProM still contains semantically unique prompts, and we denote it as VidProS. More semantically unique prompts imply covering a broader range of topics, increasing the diversity and richness of the content.
4 Comparison with DiffusionDB
In this section, we provide a comparison of the basic information between our VidProM and DiffusionDB [22], which is a large-scale prompt dataset for image generation. As shown in Table 1, the comparison is conducted in three aspects: prompts, images or videos, and licenses. We have the following observations:
Prompt. (1) Although the total number of unique prompts in our VidProM and DiffusionDB are similar, VidProM contains significantly more () semantically unique prompts. This shows VidProM is a more diverse and representative dataset. (2) Unlike DiffusionDB, which uses the OpenAI-CLIP embedding method, our approach leverages the latest OpenAI text-embedding model, namely text-embedding-3-large. One advantage of this approach is its ability to accept much longer prompts compared to CLIP, supporting up to 8192 tokens versus CLIP’s 77 tokens. As illustrated by the comparison of the number of words per prompt between VidProM and DiffusionDB in Fig. 4, the prompts used for generating videos are much more longer. Therefore, the capability of text-embedding-3-large is particularly suitable for them. Another advantage is text-embedding-3-large has stronger performance than CLIP on several standard benchmarks, potentially benefiting users of our VidProM. (3) The time span for collecting prompts in VidProM is much longer than in DiffusionDB. We collect prompts written by real users over a period of months, while DiffusionDB’s collection spans only month. A longer collection period implies a broader range of topics and themes covered, as demonstrated by the comparison of the number of semantically unique prompts.
Images or videos. DiffusionDB focuses on images, while our VidProM is specialized in videos. Therefore, given that generating videos is much more expensive than images, it is reasonable that the number of videos in VidProM is smaller than the number of images in DiffusionDB. We also make several efforts to mitigate this disadvantage: (1) The number of source diffusion models for our VidProM is much larger than those of DiffusionDB. Our videos are generated by state-of-the-art text-to-video diffusion models, while DiffusionDB contains only images generated by Stable Diffusion. As a result, the average repetition rate per source is only for our VidProM compared to about for DiffusionDB. (2) We devote significantly more resources to VidProM. Unlike DiffusionDB, which only collects images through web scraping, we also deploy three open-source text-to-video models on our local servers, dedicating over V100 GPU hours to video generation. These efforts result in our VidProM containing more than million seconds of video.
Licenses. Unlike DiffusionDB, which adheres to the CC0 1.0 license from the Stable Diffusion Discord, the prompts and videos generated by Pika in our VidProM are licensed under the CC BY-NC 4.0 license from the Pika Discord, allowing for non-commercial use with attribution. Additionally, similar to their original repositories, the videos from VideoCraft2, Text2Video-Zero, and ModelScope are released under the Apache license, the CreativeML Open RAIL-M license, and the CC BY-NC 4.0 license, respectively.
Aspects |
Details |
DiffusionDB [22] |
VidProM |
| Prompts |
No. of unique prompts |
No. of semantically unique prompts |
Embedding of prompts |
OpenAI-CLIP |
OpenAI-text-embedding-3-large |
Maximum length of prompts |
77 tokens |
8192 tokens |
Time span |
Aug 2022 |
Jul 2023 Feb 2024 |
| Images |
No. of images/videos |
million images |
million videos |
| or |
No. of sources |
| Videos |
Average repetition rate per source |
Collection method |
Web scraping |
Web scraping + Local generation |
GPU consumption |
V100 GPU hours |
Total seconds |
seconds |
Legal |
Open source license |
CC0 1.0 |
CC BY-NC 4.0 and others |
5 Prompt Analysis
In this section, we analyze the prompts in our VidProM dataset from two perspectives: (1) the necessity of introducing a new prompt dataset and (2) people’s preferences in video generation.
Why do we need a new prompt dataset for text-to-video generation? As there exists a prompt dataset (DiffusionDB) for text-to-image generation, we provide an in-depth analysis on how the prompts in our VidProM differ from those in DiffusionDB.
Firstly, as shown in Fig. 4 (a), the semantics of the prompts differ in three aspects: (1) Time dimension: text-to-video prompts usually need to include a description of the time dimension, such as ‘changes in actions’ and ‘transitions in scenes’; while text-to-image prompts typically describe a scene or object. (2) Dynamic description: text-to-video prompts often need to describe the dynamic behavior of objects, such as ‘flying’, ‘working’, and ‘writing’; while text-to-image prompts focus more on describing the static appearance of objects. (3) Duration: text-to-video prompts may need to specify the duration of the video or an action, such as ‘a long time’ and ‘1-minute’, while text-to-image prompts do not need to consider the time factor.
Secondly, as shown in Fig. 4 (a), text-to-video prompts are generally more complex and longer than text-to-image prompts, due to the need to describe additional dimensions and dynamic changes. This phenomenon is also observed in the prompts used by Sora. For instance, the prompt for the ‘Tokyo Girl’ video contains words, while the longest prompt on the OpenAI official website comprises words. As depicted in Fig. 4 (b), our VidProM dataset prominently features this characteristic: (1) the number of prompts with more than words is nearly for our VidProM, compared to only about for DiffusionDB; and (2) our VidProM still has over prompts with more than words, whereas this number is close to for DiffusionDB.
Finally, as shown in Fig. 4 (c), the prompts from our VidProM dataset and DiffusionDB exhibit different distributions. We use the text-embedding-3-large model to re-extract the features of the prompts in DiffusionDB and randomly select prompt features from both our VidProM dataset and DiffusionDB, respectively. By visualizing with t-SNE [25], we find that these prompts have significantly different distributions and are nearly linearly separable.
What kind of videos are people interested in generating? We use two methods, i.e. WordCloud and WizMap [26], to analyze the text-to-video prompts preferred by people. WordCloud is a visual representation of text, in which the size of each word indicates its frequency of occurrence within the text. WizMap [26] is a visualization tool to navigate and interpret large-scale embedding spaces with ease. From the visualization results shown in Fig. 5, we conclude that:
(1) The popular words include modern, motion, close, forest, sky, and so on. In Fig. 5 (a), we can identify the most frequent words, which can be roughly divided into two groups: words describing the characteristics of the generated videos, indicating the desired style, such as ‘modern’, ‘motion’, and ‘close’; and words describing the subject of the generation, such as ‘forest’ and ‘sky’.
(2) The preferred topics include humans, science fiction, animals, and others. Firstly, the topic of humans is in the lower and middle parts of Fig. 5 (b): the middle part depicts general behaviors of men and women, such as walking, while the lower part showcases famous individuals like Joe Biden, Trump, and Jackson. Secondly, the science fiction topic is located in the upper part, including dragons flying in the sky from Chinese tradition and aliens with spaceships from the West. Thirdly, people also like cute animals, such as cats and dogs (in the lower right), as well as wild animals, such as tigers and lions (in the middle right). Finally, other topics, such as cars (on the right) and theology (in the lower left), are also mentioned in the prompts.
6 Inspiring New Research
The proposed million-scale VidProM dataset inspires new directions for researchers to (1) develop better, more efficient, and safer text-to-video diffusion models, and (2) advance multimodal learning from synthetic videos.
Video Generative Model Evaluation aims to assess the performance and quality of text-to-video generative models. Current evaluation efforts, such as [42, 43, 44, 45], are conducted using carefully designed and small-scale prompts. Our VidProM dataset brings imagination to this field: (1) Instead of using carefully designed prompts, researchers could consider whether their models can generalize to prompts from real users, which would make their evaluation more practical. (2) Performing evaluations on large-scale datasets will make their arguments more convincing and insightful.
Text-to-Video Diffusion Model Development aims to create diffusion models capable of converting textual descriptions into dynamic and realistic videos. The current methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] are trained on caption-(real)-video pairs. Two natural questions arise: (1) Will the domain gap between captions and prompts from real users hinder these models’ performance? (2) Can researchers train or distill new text-to-video diffusion models on prompt-(generated)-video pairs? The studies of training text-to-video diffusion models on our VidProM may provide answers to these two questions.
Text-to-Video Prompt Engineering is to optimize the interaction between humans and text-to-video models, ensuring that the models understand the task at hand and generate relevant, accurate, and coherent videos. The prompt engineering field has gained attention in large language models [52, 53], text-to-image diffusion models [54, 55], and visual in-context learning [56, 57]. However, as far as we know, there is no related research in the text-to-video community. Our VidProM provides an abundant resource for text-to-video prompt engineering.
Efficient Video Generation. The current text-to-video diffusion models are very time-consuming. For example, on a single V100 GPU, ModelScope [5] requires seconds, while VideoCrafter2 [3] needs seconds to generate a video, respectively. Our large-scale VidProM provides a unique opportunity for efficient video generation. Given an input prompt, a straightforward approach is to search for the most closely related prompts in our VidProM and reconstruct a video from the corresponding existing videos, instead of generating a new video from noise.
Fake Video Detection aims to distinguish between real videos and those generated by diffusion models. While there are some works [27, 28, 29, 30, 31, 32, 33, 34, 35, 36] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames. For instance, a model trained on images generated by Stable Diffusion [37] may fail to identify frames from videos generated by Pika [9] or Sora [6] as fake. (2) The efficiency problem: Currently, there is no detector that can take an entire video as input. As a result, we must use fake image detectors to examine all or representative frames, which can be time-consuming. With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from different diffusion models to further improve the models’ performance.
Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of their training sources. Videos generated with replicated content may infringe on the copyrights of the original videos. However, current video copy detection researches [38, 39, 40, 41] focus on detecting hand-crafted copies generated by transformations like horizontal flips and random crops, and overlook the challenge from diffusion models. Researchers can utilize our dataset VidProM as input queries and employ existing video copy detection models to pre-filter potential replications. Subsequently, these filtered videos can be manually annotated to determine whether they are indeed replications and regarded as training sources.
Multimodal Learning from Synthetic Videos. Multimodal learning related to text-video interactions include tasks such as video-text retrieval [46, 47, 48] and video captioning [49, 50, 51]. They are typically trained on large-scale video datasets collected from online sources such as YouTube and Shutterstock. However, there are two main concerns: (1) copyright issues, for example, the WebVid-10M dataset [15] is no longer publicly available due to its infringement on Shutterstock’s copyright; and (2) the difficulty in collecting high-quality video-text data, as manual labeling is time-consuming. Our VidProM offers a promising research direction, i.e. learning from synthetic videos, which are free from copyright issues and do not require any manual labeling.
7 Conclusion
This paper introduces VidProM, the first dataset comprising 1.67 million unique text-to-video prompts, 6.69 million videos generated by four state-of-the-art diffusion models, along with NSFW probabilities, 3072-dimensional prompt embeddings, and additional related metadata. We demonstrate the time-consuming and expensive process of curating the dataset and provide a comparison between it and a text-to-image prompt dataset, DiffusionDB. Based on our analysis of the prompts, we find that (1) there is a need for a specialized dataset containing text-to-video prompts, and (2) people show interest in topics such as humans, science fiction, and animals when generating videos. Finally, we outline the potential research directions inspired by our VidProM and hope it will advance research in the text-video domain. In the future, we plan to enrich our dataset by incorporating high-quality videos generated by Sora with our long and detailed prompts.

### v2

@@
 The arrival of Sora marks a new era for text-to-video diffusion models, bringing significant advancements in video generation and potential applications.
-However, Sora, as well as other text-to-video diffusion models, highly relies on the prompts, and there is no publicly available dataset featuring a study of text-to-video prompts.
+However, Sora, along with other text-to-video diffusion models, is highly reliant on prompts, and there is no publicly available dataset that features a study of text-to-video prompts.
 In this paper, we introduce VidProM, the first large-scale dataset comprising 1.67 Million unique text-to-Video Prompts from real users.
-Additionally, the dataset includes 6.69 million videos generated by four state-of-the-art diffusion models and some related data.
-We initially demonstrate the curation of this large-scale dataset, which is a time-consuming and costly process.
-Subsequently, we show how the proposed VidProM differs from DiffusionDB, a large-scale prompt-gallery dataset for image generation.
-Based on the analysis of these prompts, we identify the necessity for a new prompt dataset specifically designed for text-to-video generation and gain insights into the preferences of real users when creating videos.
-Our large-scale and diverse dataset also inspires many exciting new research areas.
-For instance, to develop better, more efficient, and safer text-to-video diffusion models, we suggest exploring text-to-video prompt engineering, efficient video generation, and video copy detection for diffusion models.
-We make the collected dataset VidProM publicly available at GitHub and Hugging Face under the CC-BY-NC 4.0 License.
+Additionally, this dataset includes 6.69 million videos generated by four state-of-the-art diffusion models, alongside some related data.
+We initially discuss the curation of this large-scale dataset, a process that is both time-consuming and costly.
+Subsequently, we underscore the need for a new prompt dataset specifically designed for text-to-video generation by illustrating how VidProM differs from DiffusionDB, a large-scale prompt-gallery dataset for image generation.
+Our extensive and diverse dataset also opens up many exciting new research areas.
+For instance, we suggest exploring text-to-video prompt engineering, efficient video generation, and video copy detection for diffusion models to develop better, more efficient, and safer models.
+The project (including the collected dataset VidProM and related code) is publicly available at https://vidprom.github.io under the CC-BY-NC 4.0 License.
 1 Introduction
@@
 However, Sora [1] and other text-to-video diffusion models [2, 3, 4, 5] heavily relies on the prompts used.
-Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which hinders the development and evaluation of these models.
+Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which hinders the development of these models and related researches.
 In this paper, we present the first systematic research on the text-to-video prompts.
-Specifically, our efforts primarily focus on building the first text-to-video prompt-gallery dataset VidProM, conducting an in-depth comparison with DiffusionDB [22], analyzing prefered topics by users, and introducing new research directions based on our VidProM.
+Specifically, our efforts primarily focus on building the first text-to-video prompt-gallery dataset VidProM, analyzing the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models, and introducing new research directions based on our VidProM.
 The demonstration of VidProM is shown in Fig.
@@
 In addition to the main dataset, we introduce a subset named VidProS, which consists of semantically unique prompts.
-In this subset, the cosine similarity between any two prompts is less than 0.8, ensuring a high level of semantic diversity.
-In-depth comparison with DiffusionDB and people preference analysis.
+That means, in this subset, the cosine similarity between any two prompts is less than 0.8, ensuring a high level of semantic diversity.
+The necessity of collecting a new prompt dataset specialized for text-to-video diffusion models.
 We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [6].
-By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [22] lies in: (1) Our VidProM contains more semantically unique prompts, which are embedded by a more advanced model and collected over a longer time span.
+By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [6] mainly lies in: (1) The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer.
+(2) Our VidProM contains more semantically unique prompts, which are embedded by a more advanced model and collected over a longer time span.
 (3) We collect videos through both web scraping and local generation, whereas DiffusionDB [6] only contains web-scraped images, resulting in our dataset being more time-consuming and costly to assemble.
-(3) The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer.
-These differences highlight the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models.
-Based on our analysis of the new prompt dataset, we conclude that (1) frequently occurring words include ‘modern’, ‘motion’, ‘close’, ‘forest’, and ‘sky’; and (2) preferred topics encompass humans, science fiction, and animals.
 Inspiring new research directions.
-The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up numerous exciting research directions.
-Researchers are focused on developing better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering.
+The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up many exciting research directions.
+With the help of our VidProM, researchers can develop better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering.
 (2) For more efficient models, researchers can search for related prompts in our VidProM and reconstruct new videos from similar existing videos, thereby avoiding the need to generate videos from scratch.
 (3) For safer models, researchers can develop specialized models to distinguish generated videos from real videos to combat misinformation, and train video copy detection models to identify potential copyright issues.
-Apart from diffusion models, text-video pairs are also utilized in multimodal learning tasks such as video-text retrieval and video captioning.
-Our prompts and synthetic videos can help mitigate copyright concerns associated with using online videos and alleviate the difficulty of collecting high-quality video-text data.
+Apart from diffusion models, text-video pairs can also be utilized in multimodal learning tasks such as video-text retrieval and video captioning.
+Our prompts and synthetic videos can help mitigate copyright concerns associated with using online videos and alleviate the difficulty of collecting video-text data.
 To sum up, this paper makes the following contributions: (1) We contribute the first text-to-video prompt-gallery dataset, VidProM, which includes million unique prompts from real users and million generated videos by state-of-the-art diffusion models.
-(2) We provide a detailed and in-depth comparison with the text-to-image prompt-gallery dataset, DiffusionDB, and highlight the necessity of VidProM as well as real users’ preference.
+(2) We highlight the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models by comparing our VidProM with DiffusionDB.
 (3) We reveal several exciting research directions inspired by VidProM and position it as a rich database for future studies.
@@
 We subsequently filter out prompts used for image-to-video generation (because the images are not publicly available) and prompts without associated videos (these prompts may have been banned by Pika or hidden by the users).
-Finally, we remove duplicate prompts and assign a UUID to each one, resulting in a total of unique prompts.
+Finally, we remove duplicate prompts and assign a UUID to each prompt, resulting in a total of unique prompts.
 Because the text-to-video prompts are significantly complex and long, we use OpenAI’s text-embedding-3-large API, which supports up to tokens, to embed all of our prompts.
-We retain the original 3072-dimensional output of the text-embedding-3-large model, allowing users of our dataset to freely reduce their dimensionality according to their preferences using Principal Component Analysis (PCA).
+We retain the original 3072-dimensional output of the text-embedding-3-large model, allowing people to freely reduce their dimensionality according to their preferences using Principal Component Analysis (PCA).
 Assigning NSFW Probabilities.
@@
 This process demands significant computational resources: we distribute the text-to-video generation across 10 servers, each equipped with 8 Nvidia V100 GPUs.
-It costs us approximately 50,631 GPU hours and results in 6.69 million videos (), totaling 24,414,747.8 seconds in duration.
-The breakdown of video lengths is as follows: 3.0 seconds for Pika, 2.0 seconds for Text2Video-Zero, 1.6 seconds for VideoCraft2, and 8.0 seconds for ModelScope.
+It costs us approximately 50,631 GPU hours and results in 6.69 million videos (), totaling 14,381,289.8 seconds in duration.
+The breakdown of video lengths is as follows: 3.0 seconds for Pika, 1.6 seconds for VideoCraft2, 2.0 seconds for Text2Video-Zero, and 2.0 seconds for ModelScope.
 Selecting Semantically Unique Prompts.
@@
 We define a dataset as containing only semantically unique prompts if, for any two arbitrary prompts, their cosine similarity calculated using text-embedding-3-large embeddings is less than .
-After semantic de-duplication, our VidProM still contains semantically unique prompts, and we denote it as VidProS.
+After semantic de-duplication (see a detailed description of this process in the Appendix (Section A)), our VidProM still contains semantically unique prompts, and we denote it as VidProS.
 More semantically unique prompts imply covering a broader range of topics, increasing the diversity and richness of the content.
-4 Comparison with DiffusionDB
-In this section, we provide a comparison of the basic information between our VidProM and DiffusionDB [22], which is a large-scale prompt dataset for image generation.
-As shown in Table 1, the comparison is conducted in three aspects: prompts, images or videos, and licenses.
-We have the following observations:
+| Aspects | Details | DiffusionDB [6] | VidProM |
+| Prompts | No. of unique prompts | ||
+| No. of semantically unique prompts | |||
+| Embedding of prompts | OpenAI-CLIP | OpenAI-text-embedding-3-large | |
+| Maximum length of prompts | 77 tokens | 8192 tokens | |
+| Time span | Aug 2022 | Jul 2023 Feb 2024 | |
+| Images | No. of images/videos | million images | million videos |
+| or | No. of sources | ||
+| Videos | Average repetition rate per source | ||
+| Collection method | Web scraping | Web scraping + Local generation | |
+| GPU consumption | - | V100 GPU hours | |
+| Total seconds | - | seconds |
+4 The Necessity of Introducing VidProM
+We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [6].
+This section highlights how our VidProM is different from this dataset from two aspects, i.e.
+basic information and semantics of prompts.
+4.1 Basic Information
+In Table 1, we provide a comparison of the basic information between our VidProM and DiffusionDB [6]. We have the following observations:
 Prompt.
@@
 These efforts result in our VidProM containing more than million seconds of video.
-Licenses.
-Unlike DiffusionDB, which adheres to the CC0 1.0 license from the Stable Diffusion Discord, the prompts and videos generated by Pika in our VidProM are licensed under the CC BY-NC 4.0 license from the Pika Discord, allowing for non-commercial use with attribution.
-Additionally, similar to their original repositories, the videos from VideoCraft2, Text2Video-Zero, and ModelScope are released under the Apache license, the CreativeML Open RAIL-M license, and the CC BY-NC 4.0 license, respectively.
-Aspects |
-Details |
-DiffusionDB [22] |
-VidProM |
-| Prompts |
-No. of unique prompts |
-No. of semantically unique prompts |
-Embedding of prompts |
-OpenAI-CLIP |
-OpenAI-text-embedding-3-large |
-Maximum length of prompts |
-77 tokens |
-8192 tokens |
-Time span |
-Aug 2022 |
-Jul 2023 Feb 2024 |
-| Images |
-No. of images/videos |
-million images |
-million videos |
-| or |
-No. of sources |
-| Videos |
-Average repetition rate per source |
-Collection method |
-Web scraping |
-Web scraping + Local generation |
-GPU consumption |
-V100 GPU hours |
-Total seconds |
-seconds |
-Legal |
-Open source license |
-CC0 1.0 |
-CC BY-NC 4.0 and others |
-5 Prompt Analysis
-In this section, we analyze the prompts in our VidProM dataset from two perspectives: (1) the necessity of introducing a new prompt dataset and (2) people’s preferences in video generation.
-Why do we need a new prompt dataset for text-to-video generation?
-As there exists a prompt dataset (DiffusionDB) for text-to-image generation, we provide an in-depth analysis on how the prompts in our VidProM differ from those in DiffusionDB.
+4.2 Semantics of prompts
+In this section, we analyze how the semantics of prompts in our VidProM dataset are different from DiffusionDB [6].
 Firstly, as shown in Fig.
@@
 Secondly, as shown in Fig.
-4 (a), text-to-video prompts are generally more complex and longer than text-to-image prompts, due to the need to describe additional dimensions and dynamic changes.
+4 (a) and (b), text-to-video prompts are generally more complex and longer than text-to-image prompts, due to the need to describe additional dimensions and dynamic changes.
 This phenomenon is also observed in the prompts used by Sora.
 For instance, the prompt for the ‘Tokyo Girl’ video contains words, while the longest prompt on the OpenAI official website comprises words.
-As depicted in Fig.
-4 (b), our VidProM dataset prominently features this characteristic: (1) the number of prompts with more than words is nearly for our VidProM, compared to only about for DiffusionDB; and (2) our VidProM still has over prompts with more than words, whereas this number is close to for DiffusionDB.
+Our VidProM dataset prominently features this characteristic: (1) the number of prompts with more than words is nearly for our VidProM, compared to only about for DiffusionDB; and (2) our VidProM still has over prompts with more than words, whereas this number is close to for DiffusionDB.
 Finally, as shown in Fig.
-4 (c), the prompts from our VidProM dataset and DiffusionDB exhibit different distributions.
-We use the text-embedding-3-large model to re-extract the features of the prompts in DiffusionDB and randomly select prompt features from both our VidProM dataset and DiffusionDB, respectively.
-By visualizing with t-SNE [25], we find that these prompts have significantly different distributions and are nearly linearly separable.
-What kind of videos are people interested in generating?
-We use two methods, i.e.
-WordCloud and WizMap [26], to analyze the text-to-video prompts preferred by people.
-WordCloud is a visual representation of text, in which the size of each word indicates its frequency of occurrence within the text.
+4 (c) and Fig.
+5, the prompts from our VidProM dataset and DiffusionDB exhibit different distributions.
+We use the text-embedding-3-large model to re-extract the features of the prompts in DiffusionDB.
+For visualizing with t-SNE [25], we randomly select prompt features from both our VidProM dataset and DiffusionDB, respectively.
+We find that these prompts have significantly different distributions and are nearly linearly separable.
+Beyond the traditional t-SNE [25], we use a recent method, WizMap [26], to analyze the topics preferred by people using all prompts from our VidProM and DiffusionDB.
 WizMap [26] is a visualization tool to navigate and interpret large-scale embedding spaces with ease.
 From the visualization results shown in Fig.
-5, we conclude that:
-(1) The popular words include modern, motion, close, forest, sky, and so on.
-In Fig.
-5 (a), we can identify the most frequent words, which can be roughly divided into two groups: words describing the characteristics of the generated videos, indicating the desired style, such as ‘modern’, ‘motion’, and ‘close’; and words describing the subject of the generation, such as ‘forest’ and ‘sky’.
-(2) The preferred topics include humans, science fiction, animals, and others.
-Firstly, the topic of humans is in the lower and middle parts of Fig.
-5 (b): the middle part depicts general behaviors of men and women, such as walking, while the lower part showcases famous individuals like Joe Biden, Trump, and Jackson.
-Secondly, the science fiction topic is located in the upper part, including dragons flying in the sky from Chinese tradition and aliens with spaceships from the West.
-Thirdly, people also like cute animals, such as cats and dogs (in the lower right), as well as wild animals, such as tigers and lions (in the middle right).
-Finally, other topics, such as cars (on the right) and theology (in the lower left), are also mentioned in the prompts.
+5, we conclude that, despite some overlaps, there are distinct differences in their interests.
+For example, both groups are interested in topics like cars, food, and cute animals.
+However, text-to-image users tend to focus on generating art paintings, whereas text-to-video users show little interest in this area.
+Instead, text-to-video users are more inclined to create general human activities, such as walking, which are rarely produced by text-to-image users.
 5 Inspiring New Research
@@
 The current methods [3, 7, 4, 8, 5, 1, 9, 10, 2, 11, 12] are trained on caption-(real)-video pairs.
-Two natural questions arise: (1) Will the domain gap between captions and prompts from real users hinder these models’ performance?
+Two natural questions arise: (1) Will the domain gap between captions and prompts from real users (see the evidence of existing domain gap in the Appendix (Section B)) hinder these models’ performance?
+For instance, of the training data for Open-Sora-Plan v1.0.0 [31] consist of landscape videos, leading to the trained model’s suboptimal performance in generating scenes involving humans and cute animals.
 (2) Can researchers train or distill new text-to-video diffusion models on prompt-(generated)-video pairs?
-The studies of training text-to-video diffusion models on our VidProM may provide answers to these two questions.
+Although some studies indicate that this approach may lead to irreversible defects [32], as we will likely run out of data on websites and some methods are proposed to prevent such collapse [33], exploring synthetic data remains a valuable endeavor.
+The studies of training text-to-video diffusion models with our VidProM may provide answers to these two questions.
 Text-to-Video Prompt Engineering is to optimize the interaction between humans and text-to-video models, ensuring that the models understand the task at hand and generate relevant, accurate, and coherent videos.
@@
 Our VidProM provides an abundant resource for text-to-video prompt engineering.
+In the Appendix (Section C), we train a large language model on our VidProM for automatic text-to-video prompt completion.
 Efficient Video Generation.
@@
 Fake Video Detection aims to distinguish between real videos and those generated by diffusion models.
-While there are some works [27, 28, 29, 30, 31, 32, 33, 34, 35, 36] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames.
+While there are some works [40, 41, 42, 43, 44, 45, 46, 47, 48, 49] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames (see experiments in the Appendix (Section D)).
 For instance, a model trained on images generated by Stable Diffusion [50] may fail to identify frames from videos generated by Pika [2] or Sora [1] as fake.
 (2) The efficiency problem: Currently, there is no detector that can take an entire video as input.
-As a result, we must use fake image detectors to examine all or representative frames, which can be time-consuming.
-With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from different diffusion models to further improve the models’ performance.
-Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of their training sources.
+As a result, to achieve higher accuracy, we may use fake image detectors to examine all or representative frames, which can be time-consuming.
+With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from more diffusion models to further improve the detection performance.
+Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of existing ones (see a detailed discussion in the Appendix (Section E)).
 Videos generated with replicated content may infringe on the copyrights of the original videos.
@@
 6 Conclusion
-This paper introduces VidProM, the first dataset comprising 1.67 million unique text-to-video prompts, 6.69 million videos generated by four state-of-the-art diffusion models, along with NSFW probabilities, 3072-dimensional prompt embeddings, and additional related metadata.
-We demonstrate the time-consuming and expensive process of curating the dataset and provide a comparison between it and a text-to-image prompt dataset, DiffusionDB.
-Based on our analysis of the prompts, we find that (1) there is a need for a specialized dataset containing text-to-video prompts, and (2) people show interest in topics such as humans, science fiction, and animals when generating videos.
-Finally, we outline the potential research directions inspired by our VidProM and hope it will advance research in the text-video domain.
-In the future, we plan to enrich our dataset by incorporating high-quality videos generated by Sora with our long and detailed prompts.
+This paper provides the first systematic research on text-to-video prompts.
+Specifically, we introduce VidProM, the first dataset comprising 1.67 million unique text-to-video prompts, 6.69 million videos generated by four state-of-the-art diffusion models, along with NSFW probabilities, 3072-dimensional prompt embeddings, and additional related metadata.
+To highlight the necessity of VidProM, we compare the basic information and the semantics of prompts of our VidProM to DiffusionDB, a text-to-image prompt dataset.
+Finally, we outline the potential research directions inspired by our VidProM, such as text-to-video prompt engineering, fake video detection, and video copy detection for diffusion models.
+We hope the curated large and diverse prompt-video dataset will advance research in the text-video domain.
+Limitation.
+We recognize that the videos currently generated are short and not of the highest quality.
+In the future, we intend to enhance our dataset by incorporating high-quality videos produced by more advanced models, like Sora, using our long and detailed prompts.
+Appendix A Semantic De-duplication Algorithm
+This section shows how we select semantically unique prompts, i.e., for any two arbitrary prompts, their cosine similarity is less than . We first show our algorithm:
+Then we analyze the efficiency of the algorithm.
+The most time-consuming part of our algorithm lies in building , which includes calculating the similarity matrix and selecting pairs with a similarity greater than .
+To complete this process with and , we distribute the workload across 8 A100 GPUs and 128 CPU cores using Faiss [61].
+It only takes approximately hours.
+Appendix B Domain Gap between Video Captions and Text-to-Video Prompts
+This section highlights the domain gap between training video captions and real user prompts.
+The domain gap may hinder the model’s ability to generate satisfactory videos due to a lack of exposure to topics of interest to users and the style difference between captions and prompts.
+In Fig.
+6, we randomly select million captions from Panda-70M [18] and visualize them with our VidProM.
+We find that their distributions are significantly different.
+On one hand, the covered topics are different.
+For instance, users may want to generate videos of city nights or superheroes like Spider-Man.
+However, in the training data, there are few or no related videos on these topics.
+One possible solution is to filter out training videos that cover topics of human interest before training text-to-video models.
+On the other hand, there may be a difference in styles of describing videos.
+For instance, users may start a prompt with “Generate a video of” and end it with “4K”.
+However, these words may not appear in the training captions.
+One suggestion is to train a text style transfer model [62, 63] to convert the style of the captions to that of the user prompts or reverse.
+Appendix C Automatic Text-to-video Prompt Completion
+This section describes the training and output of our text-to-video prompt completion model.
+Training details.
+Similar to text2image-prompt-generator [64], we treat the automatic text-to-video prompt completion as a general text generation task.
+The training code is available at: https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_clm.py.
+We fine-tune the recent state-of-the-art large language model, Mistral-7B-v0.1 [65], on our dataset VidProM.
+The fine-tuning process is completed using DeepSpeed [66] with 8 A100 GPUs in 2 hours.
+Demonstration.
+We release our trained model on: https://huggingface.co/WenhaoWang/AutoT2VPrompt.
+It can be freely downloaded and tested under CC BY-NC 4.0 License, and we also demonstrate some of its generation across various topics below.
+Appendix D Benchmarking Existing Fake Image Detection on VidProM
+This section shows the current fake image detection methods fail to generalize to fake video detection.
+Implementation details.
+To benchmark existing fake image detection algorithms on our dataset, we test eight open-source, state-of-the-art models [42, 44, 43, 67, 46, 45, 47, 49] available at https://github.com/Ekko-zn/AIGCDetectBenchmark.
+We randomly select 10,000 videos generated by Pika [2], VideoCraft2 [4], Text2Video-Zero [3], and ModelScope [5], respectively, to serve as fake samples.
+Similarly, we choose 10,000 real videos randomly from DVSC2023 [68].
+Since none of the state-of-the-art models can process entire videos directly, we extracted the middle frame from each video to use as the input image for each model.
+We evaluate the models using two metrics: Accuracy and Mean Average Precision (mAP).
+Experimental results.
+The performance of the models is detailed in Tables 2 and 3.
+Our observations reveal that all existing models struggle to differentiate between fake and real videos: (1) Methods initially designed for detecting GAN-generated images, such as [42] and [46], are ineffective with videos produced by diffusion models, likely because they are trained on images from significantly different generative methods.
+(2) Techniques specifically developed for the diffusion process [47] or for generalizing across various generative models [49] also fail in detecting fake videos, indicating that the key features identified for spotting diffusion-generated images may not be as widely applicable as previously assumed.
+(3) Interestingly, methods that rely on traditional cues, such as global texture statistics [67] and frequency analysis [44], are the most effective.
+This underscores the continuing importance of traditional image processing knowledge.
+| Accuracy () | Pika | VideoCraft2 | Text2Video-Zero | ModelScope | Average |
+| CNNSpot [42] | 51.17 | 50.18 | 49.97 | 50.31 | 50.41 |
+| FreDect [44] | 50.07 | 54.03 | 69.88 | 69.94 | 60.98 |
+| Fusing [43] | 50.60 | 50.07 | 49.81 | 51.28 | 50.44 |
+| Gram-Net [67] | 84.19 | 67.42 | 52.48 | 50.46 | 63.64 |
+| LGrad [46] | 53.73 | 51.75 | 41.05 | 60.22 | 51.69 |
+| LNP [45] | 43.48 | 45.10 | 47.50 | 45.21 | 45.32 |
+| DIRE [47] | 50.53 | 49.95 | 48.96 | 48.32 | 49.44 |
+| UnivFD [49] | 49.41 | 48.65 | 49.58 | 57.43 | 51.27 |
+| mAP () | Pika | VideoCraft2 | Text2Video-Zero | ModelScope | Average |
+| CNNSpot [42] | 54.63 | 41.12 | 44.56 | 46.95 | 46.82 |
+| FreDect [44] | 47.82 | 56.67 | 75.31 | 64.15 | 60.99 |
+| Fusing [43] | 57.64 | 41.64 | 40.51 | 56.09 | 48.97 |
+| Gram-Net [67] | 94.32 | 80.72 | 57.73 | 43.54 | 69.08 |
+| LGrad [46] | 54.49 | 53.21 | 36.69 | 66.53 | 52.73 |
+| LNP [45] | 44.28 | 44.08 | 46.81 | 39.62 | 43.70 |
+| DIRE [47] | 49.21 | 50.44 | 44.52 | 48.64 | 48.20 |
+| UnivFD [49] | 48.63 | 42.36 | 48.46 | 70.75 | 52.55 |
+Appendix E Video Copy Detection for Diffusion Models
+This sections shows whether the videos generated by text-to-video diffusion models replicate content of their training or existing videos.
+E.1 Experimental Setup
+The original videos for matching.
+We randomly select million videos from the training source of VideoCraft2 [4] and ModelScope [5] to analyze the extent to which generated videos replicate their training data.
+We acknowledge that using this source may result in an underestimation of the replication ratio, but it is sufficient for analyzing the phenomenon of replication.
+E.2 Observations
+Qualitative observations.
+In Fig.
+7, we demonstrate that text-to-video diffusion models can replicate the content from their training data or existing videos.
+This replication is acceptable for fair use purposes such as education, news reporting, and parody.
+However, misusing videos that contain replicated content with copyright could constitute an infringement.
+For example, in Fig.
+7, the video generated by the commercial model Pika closely replicate the content of the famous painting, “The Persistence of Memory”, which is copyrighted by the Gala-Salvador Dalí Foundation.
+If someone uses this generated video for profit without permission, it could potentially constitute copyright infringement.
+Quantitative observations.
+We calculate the maximum normalized score for each generated video by comparing it to all reference videos.
+As shown in Fig.
+8, we count the number of generated videos with a maximum normalized score exceeding 0.
+Generally, a normalized score greater than 0 suggests a high likelihood of replicated content [69, 70, 68].
+Our observations include: (1) Although replication occurs, it represents only a very small proportion of the generated videos.
+For instance, only about of the videos generated by Text2Video-Zero have a maximum normalized score greater than 0.
+(2) The commercial model Pika has the lowest replication rate, while the open-source model Text2Video-Zero has the highest.
+This difference may be partly because commercial models have implemented measures to prevent replication and mitigate related legal risks.
+E.3 Limitations and Future Directions
+In Fig.
+9, we visualize some failure cases gained by the current video copy detection models.
+We observe that, although the generated videos replicate content from existing ones, the video copy detection model fails to produce a normalized score higher than 0.
+The underlying issue is that these models are trained only on manually-generated transformations (as shown in Fig.
+10), and therefore, they do not generalize well to replication patterns produced by diffusion models.
+To establish video copy detection for diffusion models, a feasible approach is to utilize existing video copy detection systems as filters to identify potential pairs containing replicated content.
+Following this, human labelers can be employed to verify these pairs.
+The validated pairs can then be used to train a new detection model.
+Appendix F Author Statement
+We, the undersigned, hereby affirm that we bear full responsibility for the content provided herein.
+We guarantee that all data and materials used and presented are compliant with applicable laws and are free from any encroachments on intellectual property rights or any other legal rights of third parties.
+In the event of any violations, we accept full liability and will address all associated legal consequences.
+Appendix G Hosting, licensing, and maintenance plan.
+The data is hosted by Hugging Face, Inc.
+at https://huggingface.co/datasets/WenhaoWang/VidProM.
+The prompts and videos generated by Pika in our VidProM are licensed under the CC BY-NC 4.0 license from the Pika Discord, allowing for non-commercial use with attribution.
+Additionally, similar to their original repositories, the videos from VideoCraft2, Text2Video-Zero, and ModelScope are released under the Apache license, the CreativeML Open RAIL-M license, and the CC BY-NC 4.0 license, respectively.
+We commit to maintaining the dataset, which includes regular updates and corrections as necessary.
+Appendix H Data Sheet for VidProM
+For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
+The arrival of Sora marks a new era for text-to-video diffusion models, bringing significant advancements in video generation and potential applications.
+However, Sora, along with other text-to-video diffusion models, is highly reliant on prompts, and there is no publicly available dataset that features a study of text-to-video prompts.
+We underscore the need for a new prompt dataset specifically designed for text-to-video generation by illustrating how VidProM differs from DiffusionDB, a large-scale prompt-gallery dataset for image generation.
+Who created the dataset (e.g., which team, research group) and on
+behalf of which entity (e.g., company, institution, organization)?
+The dataset was created by Wenhao Wang (University of Technology Sydney), Yifan Sun (Baidu Inc.), and Yi Yang (Zhejiang University).
+Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
+Funded in part by Faculty of Engineering and Information Technology Scholarship, University of Technology Sydney.
+Any other comments?
+None.
+Composition
+What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
+Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)?
+Please provide a description.
+Each instance consists of a text-to-video prompt, a UUID, a timestamp, six NSFW probabilities, a 3072-dimensional prompt embedding, and four generated videos from Pika, VideoCraft2, Text2Video-Zero, and ModelScope.
+How many instances are there in total (of each type, if appropriate)?
+There are 1,672,243 instances in total.
+Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is
+a sample, then what is the larger set? Is the sample representative of the
+larger set (e.g., geographic coverage)? If so, please describe how this
+representativeness was validated/verified.
+If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
+The dataset contains all possible instances up to February 2024.
+What data does each instance consist of? “Raw” data (e.g., unprocessed text or images)or features? In either case, please provide a description.
+Each instance consists of a text-to-video prompt, a UUID, a timestamp, six NSFW probabilities, a 3072-dimensional prompt embedding, and four generated videos from Pika, VideoCraft2, Text2Video-Zero, and ModelScope.
+Is there a label or target associated with each instance? If so, please provide a description.
+The labels associated with each video are the prompts and other related data.
+Is any information missing from individual instances?
+If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable).
+This does not include intentionally removed information, but might include, e.g., redacted text.
+Everything is included. No data is missing.
+Are relationships between individual instances made explicit (e.g.,
+users’ movie ratings, social network links)? If so, please describe
+how these relationships are made explicit.
+Not applicable.
+Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
+No. This dataset is not for ML model benchmarking. Researchers can use any subsets of it.
+Are there any errors, sources of noise, or redundancies in the
+dataset? If so, please provide a description.
+No. All videos and prompts are extracted as is from the Discord chat log.
+Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
+The dataset is entirely self-contained.
+Does the dataset contain data that might be considered confidential
+(e.g., data that is protected by legal privilege or by doctor–patient
+confidentiality, data that includes the content of individuals’ nonpublic communications)? If so, please provide a description.
+Unknown to the authors of the datasheet.
+It is possible that some prompts contain sensitive information.
+However, it would be rare, as the Pika Discord has rules against writing personal information in the prompts, and there are moderators removing messages that violate the Discord rules.
+Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
+We collect videos and their prompts from the Pika discord server.
+Even though the discord server has rules against users sharing any NSFW (not suitable for work, such as sexual and violent content) and illegal images, VidProM still contains some NSFW videos and prompts that were not removed by the server moderators.
+To mitigate this, we provide six separate NSFW probabilities, allowing researchers to determine a suitable threshold for filtering out potentially unsafe data specific to their tasks.
+Does the dataset identify any subpopulations (e.g., by age, gender)?
+If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
+No.
+Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other
+data) from the dataset? If so, please describe how.
+No.
+Any other comments?
+None.
+Collection
+How was the data associated with each instance acquired? Was
+the data directly observable (e.g., raw text, movie ratings), reported by
+subjects (e.g., survey responses), or indirectly inferred/derived from other
+data (e.g., part-of-speech tags, model-based guesses for age or language)?
+If the data was reported by subjects or indirectly inferred/derived from
+other data, was the data validated/verified? If so, please describe how.
+The data was directly observed from the Pika Discord Channel.
+It was gathered from channels where users can generate videos by interacting with a bot, which consisted of messages of user generated videos and the prompts used to generate those images.
+What mechanisms or procedures were used to collect the data
+(e.g., hardware apparatuses or sensors, manual human curation,
+software programs, software APIs)? How were these mechanisms or
+procedures validated?
+The data was gathered using a DiscordChatExporter [23], which collected videos and chat messages from each channel specified.
+We then extracted and linked prompts to videos.
+Random videos and prompts were selected and manually verified to validate the prompt-video mapping.
+If the dataset is a sample from a larger set, what was the sampling
+strategy (e.g., deterministic, probabilistic with specific sampling
+probabilities)?
+VidProM does not sample from a larger set.
+Who was involved in the data collection process (e.g., students,
+crowdworkers, contractors) and how were they compensated (e.g.,
+how much were crowdworkers paid)?
+No crowdworkers are needed in the data collection process.
+Over what timeframe was the data collected? Does this timeframe
+match the creation timeframe of the data associated with the instances
+(e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
+All messages were generated between July 2023 and February 2023.
+VidProM includes the generation timestamps of all videos.
+Were any ethical review processes conducted (e.g., by an institutional review board)?
+If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
+There were no ethical review processes conducted.
+Did you collect the data from the individuals in question directly,
+or obtain it via third parties or other sources (e.g., websites)?
+The data was directly obtained from individual messages in the Discord server.
+Were the individuals in question notified about the data collection?
+If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
+Users of the channel were not notified about this specific gathering of data but agree to open-source their input and output under the Creative Commons Noncommercial 4.0 Attribution International License (CC BY-NC 4.0).
+The exact language is as follows:
+You hereby grant Mellis and other users a license to any of your Inputs and Outputs that you make available to other users on the Service under the Creative Commons Noncommercial 4.0 Attribution International License (as accessible here: https://creativecommons.org/licenses/by-nc/4.0/legalcode).
+Did the individuals in question consent to the collection and use
+of their data? If so, please describe (or show with screenshots or other
+information) how consent was requested and provided, and provide a
+link or other access point to, or otherwise reproduce, the exact language
+to which the individuals consented.
+By using the server and tools, users consented to the regulations posed by Pika Lab, the company that both made Pika text-to-video model and runs the Discord server.
+This implies consent by using the tool.
+The exact wording is as follows:
+You hereby grant Mellis and other users a license to any of your Inputs and Outputs that you make available to other users on the Service under the Creative Commons Noncommercial 4.0 Attribution International License (as accessible here: https://creativecommons.org/licenses/by-nc/4.0/legalcode).
+If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
+If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
+Users will have the option to report harmful content or withdraw videos they created by emailing Wenhao Wang (wangwenhao0716@gmail.com).
+Has an analysis of the potential impact of the dataset and its use
+on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including
+the outcomes, as well as a link or other access point to any supporting
+documentation.
+No analysis has been conducted.
+Any other comments?
+None.
+Preprocessing
+Was any preprocessing/cleaning/labeling of the data done (e.g.,
+discretization or bucketing, tokenization, part-of-speech tagging,
+SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the
+remaining questions in this section.
+No, we provide the original prompts and generated videos.
+Was the “raw” data saved in addition to the preprocessed/cleaned/labeled
+data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
+Yes, raw data is saved.
+Is the software that was used to preprocess/clean/label the data
+available? If so, please provide a link or other access point.
+Not applicable.
+Any other comments?
+None.
+Uses
+Has the dataset been used for any tasks already? If so, please provide a description.
+No.
+Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
+No.
+What (other) tasks could the dataset be used for?
+This dataset can be used for video generative model evaluation, text-to-video diffusion model development, text-to-video prompt engineering, efficient video generation, fake video detection, video copy detection for diffusion models, and multimodal learning from synthetic videos.
+Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
+For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)?
+If so, please provide a description.
+Is there anything a dataset consumer could do to mitigate these risks or harms?
+There is minimal risk for harm: the data were already public.
+Personally identifiable data (e.g., discord usernames) were not collected.
+Are there tasks for which the dataset should not be used? If so, please provide a description.
+All tasks that utilize this dataset should follow the licensing policies and the regulations posed by Pika Lab.
+Any other comments?
+None.
+Distribution
+Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
+Yes, the dataset is publicly available on the internet.
+How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
+The dataset is distributed on the project website: https://vidprom.github.io/.
+The dataset shares the same DOI as this paper.
+When will the dataset be distributed?
+The dataset is released on March 12nd, 2024.
+Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
+If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
+The prompts and videos generated by Pika in our VidProM are licensed under the CC BY-NC 4.0 license from the Pika Discord, allowing for non-commercial use with attribution.
+Additionally, similar to their original repositories, the videos from VideoCraft2, Text2Video-Zero, and ModelScope are released under the Apache license, the CreativeML Open RAIL-M license, and the CC BY-NC 4.0 license, respectively.
+Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
+If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
+No.
+Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
+If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
+No.
+Any other comments?
+None.
+Maintenance
+Who will be supporting/hosting/maintaining the dataset?
+The authors of this paper will be supporting and maintaining the dataset.
+How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
+The contact information of the curators of the dataset is listed on the project website: https://vidprom.github.io/.
+Is there an erratum? If so, please provide a link or other access point.
+There is no erratum for our initial release.
+Errata will be documented in future releases on the dataset website.
+Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
+If so, please describe how often, by whom, and how updates will be communicated to dataset consumers (e.g., mailing list, GitHub)?
+Yes, we will monitor cases when users can report harmful images and creators can remove their videos.
+We will update the dataset bimonthly.
+Updates will be posted on the project website https://vidprom.github.io/.
+In the future, we intend to enhance our dataset by incorporating high-quality videos produced by more advanced models, like Sora, using our long and detailed prompts.
+If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)?
+If so, please describe these limits and explain how they will be enforced.
+People can write an email to remove specific instances from VidProM.
+Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
+We will continue to support older versions of the dataset.
+If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
+If so, please provide a description.
+Will these contributions be validated/verified?
+If so, please describe how.
+If not, why not?
+Is there a process for communicating/distributing these contributions to dataset consumers?
+If so, please provide a description.
+Anyone can extend/augment/build on/contribute to VidProM.
+Potential collaborators can contact the dataset authors.
+Any other comments?
+None.
+NeurIPS Paper Checklist
+1.
+Claims
+Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+Answer: [Yes]
+Justification: We follow this common practice for writing a paper.
+Guidelines:
+The answer NA means that the abstract and introduction do not include the claims made in the paper.
+The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+2.
+Limitations
+Question: Does the paper discuss the limitations of the work performed by the authors?
+Answer: [Yes]
+Justification: See the Limitation at the end of our main paper.
+Guidelines:
+The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+The authors are encouraged to create a separate "Limitations" section in their paper.
+The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+3.
+Theory Assumptions and Proofs
+Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+Answer: [N/A]
+Justification: This is not a theory paper.
+Guidelines:
+The answer NA means that the paper does not include theoretical results.
+All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+All assumptions should be clearly stated or referenced in the statement of any theorems.
+The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+Theorems and Lemmas that the proof relies upon should be properly referenced.
+4.
+Experimental Result Reproducibility
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+Answer: [Yes]
+Justification: The project (including the collected dataset VidProM and related code) is publicly available at https://vidprom.github.io under the CC-BY-NC 4.0 License.
+Guidelines:
+The answer NA means that the paper does not include experiments.
+If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+5.
+Open access to data and code
+Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+Answer: [Yes]
+Justification:The project (including the collected dataset VidProM and related code) is publicly available at https://vidprom.github.io under the CC-BY-NC 4.0 License.
+Guidelines:
+The answer NA means that paper does not include experiments requiring code.
+Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+6.
+Experimental Setting/Details
+Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+Answer: [Yes]
+Justification: The experimental settings and details are provided in the main paper as well as the Appendix.
+Guidelines:
+The answer NA means that the paper does not include experiments.
+The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+The full details can be provided either with the code, in appendix, or as supplemental material.
+7.
+Experiment Statistical Significance
+Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+Answer: [No]
+Justification: All the experimental results are significant and stable, and error bars are not reported because it would be too computationally expensive.
+Guidelines:
+The answer NA means that the paper does not include experiments.
+The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+The assumptions made should be given (e.g., Normally distributed errors).
+It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+8.
+Experiments Compute Resources
+Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+Answer: [Yes]
+Justification: We include the experiments compute resources in each related section.
+Guidelines:
+The answer NA means that the paper does not include experiments.
+The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+9.
+Code Of Ethics
+Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+Answer: [Yes]
+Justification: We follow the NeurIPS Code of Ethics.
+Guidelines:
+The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+10.
+Broader Impacts
+Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+Answer: [Yes]
+Justification: Please see the end of Conclusion and the Data Sheet.
+Guidelines:
+The answer NA means that there is no societal impact of the work performed.
+If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+11.
+Safeguards
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+Answer: [N/A]
+Justification: Our paper poses no such risks.
+Guidelines:
+The answer NA means that the paper poses no such risks.
+Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+12.
+Licenses for existing assets
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+Answer: [Yes]
+Justification: We cite all the assets used in our paper, and all the corresponding licenses are mentioned and respected.
+Guidelines:
+The answer NA means that the paper does not use existing assets.
+The authors should cite the original paper that produced the code package or dataset.
+The authors should state which version of the asset is used and, if possible, include a URL.
+The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+13.
+New Assets
+Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+Answer: [Yes]
+Justification: The project (including the collected dataset VidProM and related code) is publicly available at https://vidprom.github.io under the CC-BY-NC 4.0 License.
+Guidelines:
+The answer NA means that the paper does not release new assets.
+Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+The paper should discuss whether and how consent was obtained from people whose asset is used.
+At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+14.
+Crowdsourcing and Research with Human Subjects
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+Answer: [N/A]
+Justification: The paper does not involve crowdsourcing nor research with human subjects.
+Guidelines:
+The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+15.
+Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+Answer: [N/A]
+Justification: We do not involve research with human subjects and there is no potential risk.
+Guidelines:
+The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.